# A Researcher’s Guide to Empirical Risk Minimization

### v1

@@ 4.2 From entropy to critical radii: bounding localized complexity
where $N_{\infty}(\varepsilon,\mathcal{G})$ is the $\varepsilon$–covering number with respect to the $P$-essential supremum metric. Note that $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\mathcal{J}_{\infty}(\delta,\mathcal{G})$. Entropy integrals enjoy a useful monotonicity: $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})$ is nondecreasing, while $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})/\delta$ is nonincreasing. Here the lower limit of integration is 0, so the integral may diverge; for example, it diverges for Hölder classes when $d>2s$. While we do not pursue this here, a standard workaround is to use a truncated entropy integral that integrates from $\delta^{2}/2$ to $\delta$, as in Corollary 14.3 of Wainwright, 2019.
@@
We now present our local maximal inequalities for Rademacher complexities. Our first result bounds the localized Rademacher complexity of $\mathcal{G}$ in terms of the uniform entropy integral and is adapted from Theorem 2.1 of Van Der Vaart and Wellner (2011). It shows that, up to universal constants, a solution to the critical inequality $\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta^{2}$ can be obtained by solving the entropy-based inequality $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\sqrt{n}\,\delta^{2}$.
@@ 1. A3)
It is useful to note that $\delta\sqrt{\log(1+1/\delta)}\lesssim\sqrt{n}\,\delta^{2}$ whenever $\delta\gtrsim\{\log(en)/n\}^{1/2}$. Consequently, if $\delta_{n}$ satisfies $\mathcal{J}_{2}(\delta_{n},\mathcal{F}_{\ell})\lesssim\sqrt{n}\,\delta_{n}^{2}$, then, by Lemma 6, the expanded radius $\widetilde{\delta}_{n}:=\delta_{n}\vee\{\log(en)/n\}^{1/2}$ satisfies the critical inequality $\mathcal{J}_{2}(\widetilde{\delta}_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell}))\lesssim\sqrt{n}\,\widetilde{\delta}_{n}^{2}$, up to constants depending on $M$. Combining Theorem 3 with Lemma 5 and Lemma 6 yields the following corollary.
@@
The corollary is stated in terms of the entropy-based critical radius $\delta_{n}$ of the original ERM class $\mathcal{F}$, rather than the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$. This simplification comes at a cost: by Lemma 6, it incurs an additional $\log(eMn)/n$ term in the regret bound. This term is typically loose and can often be improved with more refined arguments (see, e.g., Theorem 14.20(b) of Wainwright, 2019 and Appendix K of Foster and Syrgkanis, 2023). In particular, Appendix B.3 gives a local concentration inequality, derived from Wainwright, 2019, that may be preferred over Theorem 2 when the loss $\ell$ satisfies A3; it allows $\log(eMn)$ to be replaced by $\log\log(Men)$. Alternatively, if one only seeks the in-probability rate $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})$, it suffices to combine Theorem 3.4.1 of Van Der Vaart and Wellner (1996) with Theorem 6 applied directly to $\mathcal{F}_{\ell}$, avoiding star-shaped hulls and the resulting extra logarithmic factors. Nonetheless, the corollary suffices for most nonparametric applications. Outside low-dimensional parametric settings—where $\delta_{n}^{2}$ can scale as $n^{-1}$ and $\log n/n$} may dominate—it yields the usual sharp rates.
4.3 Illustrative example
@@ 5.1 Weighted ERM and regret transfer
The following theorem characterizes the excess regret incurred by having to estimate the weight function. In what follows, let $f_{0}\in\arg\min_{f\in\mathcal{F}}R(f;w_{0})$ minimize the population risk and let $\hat{f}_{n}$ be any random element of $\mathcal{F}$. We define the regret $\mathrm{Reg}(f;w):=R(f;w)-\inf_{f\in\mathcal{F}}R(f;w),$ where $R(f;w):=P\{w(\cdot)\,\ell(\cdot,f)\}.$ We also impose a weighted variant of the Bernstein condition.
@@ 2. B3)
Condition B2 holds under the sufficient conditions of Lemma 1. Condition B3B3-ii ensures that $f$ and $g$ enter the loss $\ell_{g}(\cdot,f)$ in a pointwise Lipschitz manner. The product structure in Condition B3B3-i is satisfied by many losses. For example, it holds for pseudo-outcome regression losses (Rubin and van der Laan, 2006; Yang et al., 2023), such as the DR-learner (van der Laan, 2013; Luedtke and van der Laan, 2016; Kennedy, 2023), which take the form $\ell_{g}(z,f)=\{g(z)-f(z)\}^{2}$ for a pseudo-outcome function $g$. In this case, one may set $m_{1}(z,f):=f(z)$, $m_{2}(z,g):=-2g(z)$, $r_{1}(z,f):=f(z)^{2}$, and $r_{2}(z,g):=g(z)^{2}$. The condition also holds for the weighted loss in Section 5.1, $\ell_{g}(z,f):=g(z)\,\ell(z,f)$, with $m_{1}(z,f):=\ell(z,f)$ and $m_{2}(z,g):=g(z)$. More generally, our analysis extends to losses that can be written as finite sums of terms with this product structure, possibly involving multiple nuisance components, as in the general loss classes studied by van der Laan et al. (2024a).
@@ 1. B4)
    (Sup-norm relation) There exist $c_{\infty}\in(0,\infty)$ and $\beta\ \geq 1/2$ such that
@@ 2. B5)
    (PAC bound for the nuisance) For all $\eta\in(0,1)$,
@@
Condition B4 always holds with $\beta=1/2$ and $c_{\infty}:=2\sup_{f\in\mathcal{F}}\|f\|_{\infty}$, but for smoothness classes it often holds with $\beta>1/2$. For finite-dimensional linear models of dimension $p$, one may take $c_{\infty}\asymp\sqrt{p}$ and $\beta=\infty$. It holds for reproducing kernel Hilbert spaces with eigenvalue decay $\lambda_{j}\asymp j^{-2\beta}$ (Mendelson and Neeman, 2010, Lemma 5.1). The embedding also holds for signed convex hulls of suitable bases (van de Geer, 2014, Lemma 2), and for $d$-variate Hölder and Sobolev classes of order $s>d/2$ on bounded domains, where $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). Condition B5 requires that the nuisance estimator satisfy a high-level PAC guarantee; for instance, this follows from Theorem 3 when the nuisance is itself obtained via empirical risk minimization over $\mathcal{G}$, with $\varepsilon_{\mathrm{nuis}}(n,\eta)$ scaling with the critical radius of $\mathcal{G}$.
The following theorem characterizes the $L^{2}(P)$ error of $\hat{f}_{n}$ in terms of the critical radius $\delta_{n,\mathcal{F}}$ of the optimization class $\mathcal{F}$, as well as the localized Rademacher complexity $\mathfrak{R}_{n}\bigl(\mathcal{G},\varepsilon_{\mathrm{nuis}}(n,\eta)\bigr)$ of the class $\mathcal{G}$ used to learn $\hat{g}$. Its proof leverages a local concentration bound for empirical inner products of the form $\{(P_{n}-P)(fg):\ f\in\mathcal{F},\ g\in\mathcal{G}\}$ (Theorem 12 in Appendix C). We denote by $\mathcal{F}$-$\mathcal{F}$ the pairwise difference class $\{f-f^{\prime}:f,f^{\prime}\in\mathcal{F}\}$, and similarly for $\mathcal{G}-\mathcal{G}$.
@@
$\delta_{n,\mathcal{F}}\ \gtrsim\ M\sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}},\qquad\delta_{n,\mathcal{G}}\ \gtrsim\ \sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}},$
where $M:=1+\sup_{g\in\mathcal{G}}\|g\|_{\infty}+\sup_{f\in\mathcal{F}}\|f\|_{\infty}$. Then, with probability at least $1-\eta$,
@@
For simplicity, suppose that $\hat{g}$ is obtained by empirical risk minimization over $\mathcal{G}$. In this case, Theorem 3 typically yields a PAC guarantee $\varepsilon_{\mathrm{nuis}}(n,\eta)\asymp\delta_{n,\mathcal{G}}$, where $\delta_{n,\mathcal{G}}$ is the critical radius of $\mathcal{G}$. Under this specialization, Theorem 9 makes explicit how the critical radii of $\mathcal{F}$ and $\mathcal{G}$ jointly control the $L^{2}(P)$ error of $\hat{f}_{n}$. The leading term $\delta_{n,\mathcal{F}}^{2}$ matches the oracle rate one would obtain if the nuisance were known. The additional price of estimating the nuisance in-sample appears through $\{\delta_{n,\mathcal{G}}^{2}\}^{4\beta/(2\beta+1)}.$ The Hölder exponent $\beta$ in Condition B4 quantifies how effectively $L^{2}(P)$ localization translates into sup-norm control, and thus governs how strongly the nuisance terms are attenuated. More generally, the theorem highlights two regimes. In an oracle regime, if $\delta_{n,\mathcal{G}}=O(\delta_{n,\mathcal{F}}^{\,(2\beta+1)/(4\beta)}),$ then the nuisance terms are $O(\delta_{n,\mathcal{F}}^{2})$ and the overall rate matches that of an oracle that knows $g_{0}$; in particular, $\|\hat{f}_{n}-\hat{f}_{0}\|^{2}\lesssim\delta_{n,\mathcal{F}}^{2}$ up to constants. By contrast, if $\delta_{n,\eta,\mathcal{G}}$ exceeds $\delta_{n,\mathcal{F}}^{\,(2\beta+1)/(4\beta)}$, then at least one nuisance term dominates $\delta_{n,\mathcal{F}}^{2}$, and the rate is governed by the nuisance complexity.
The following corollary shows that the oracle rate can be attained under the mild Donsker-type requirement $\delta_{n,\mathcal{G}}^{2}=O(n^{-1/2})$ (and likewise $\varepsilon_{\mathrm{nuis}}^{2}(n,\eta)=O(n^{-1/2})$) for function classes satisfying the critical-radius scaling $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$.
@@
[…]
@@
It turns out that requirement (ii) holds for many standard function classes. For Hölder and Sobolev classes in dimension $d$ with smoothness $s>d/2$, one typically has (11) with $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). The same scaling also holds for a broad class of reproducing kernel Hilbert spaces (RKHS) whose kernel-operator eigenvalues satisfy $\lambda_{j}\asymp j^{-2\beta}$. In particular, the sup-norm bound in Condition B4 follows from Lemma 5.1 of Mendelson and Neeman (2010). Likewise, the local Rademacher complexity scales as $\mathfrak{R}_{n}\bigl(\delta,\mathcal{F}-\mathcal{F}\bigr)\asymp\delta^{1-\tfrac{1}{2\beta}}$, with critical radius $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$ (Wainwright, 2019).
The $n^{-1/4}$-rate condition $\varepsilon_{\mathrm{nuis}}(n,\eta)=O(n^{-1/4})$ imposed by (i), together with the scalings in (ii), also appears in the R-learner analyses of Nie and Wager (2021), which study ERM with cross-fitted nuisance estimation over RKHS. This suggests that, at least for RKHS, the requirements for oracle rates are similar with and without cross-fitting, provided that $\mathcal{G}$ satisfies a Donsker-type condition and that $\delta_{n,\mathcal{G}}\asymp\varepsilon_{\mathrm{nuis}}(n,\eta)$, as is typical when $\hat{g}$ is obtained via ERM over $\mathcal{G}$. Finally, our analysis establishes the oracle rate only up to multiplicative constants; a sharper argument, as in Theorem 5 of van der Laan et al. (2024a), also recovers the oracle constant.

### v2

@@ 4.2 From entropy to critical radii: bounding localized complexity
where $N_{\infty}(\varepsilon,\mathcal{G})$ is the $\varepsilon$–covering number with respect to the $P$-essential supremum metric. Note that $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\mathcal{J}_{\infty}(\delta,\mathcal{G})$. Entropy integrals enjoy a useful monotonicity: $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})$ is nondecreasing, while $\delta\mapsto\mathcal{J}_{2}(\delta,\mathcal{G})/\delta$ is nonincreasing. Here the lower limit of integration is 0, so the integral may diverge; for example, it diverges for Hölder classes when $d>2s$. We do not pursue this here, but a standard workaround is to use a truncated entropy integral, as in Corollary 14.3 of Wainwright, 2019 and Section 4.1 of Foster and Syrgkanis (2023), which excludes a small neighborhood of 0. The ERM analysis can then be modified to control the regret only on the event that $\hat{f}_{n}$ is not already sufficiently close to $f_{0}$.
@@
We now present local maximal inequalities for Rademacher complexities. Our first result bounds the localized Rademacher complexity of $\mathcal{G}$ in terms of the uniform entropy integral and is adapted from Theorem 2.1 of Van Der Vaart and Wellner (2011). It shows that, up to universal constants, a solution to the critical inequality $\mathfrak{R}_{n}(\mathcal{G},\delta)\leq\delta^{2}$ can be obtained by solving the entropy-based inequality $\mathcal{J}_{2}(\delta,\mathcal{G})\leq\sqrt{n}\,\delta^{2}$. Related extensions that allow for unbounded function classes or envelopes appear in Chernozhukov et al. (2014) and Vaart and Wellner (2023).
@@
Let $\delta>0$. Assume $\delta_{\infty}:=\sup_{f\in\mathcal{G}(\delta)}\|f\|_{\infty}<\infty$ Then, there exists a universal constant $C\in(0,\infty)$ such that
@@
Let $\delta>0$. Define $\mathcal{G}(\delta):=\{f\in\mathcal{G}:\|f\|\leq\delta\}$ and assume $B:=1\vee\mathcal{J}_{\infty}\left(\infty,\mathcal{G}\right)<\infty$. Then there exists a universal constant $C\in(0,\infty)$ such that
@@ 1. A3)
It is useful to note that $\delta\sqrt{\log(1+1/\delta)}\lesssim\sqrt{n}\,\delta^{2}$ whenever $\delta\gtrsim\{\log(en)/n\}^{1/2}$. Consequently, if $\delta_{n}$ satisfies $\mathcal{J}_{2}(\delta_{n},\mathcal{F}_{\ell})\lesssim\sqrt{n}\,\delta_{n}^{2}$, then, by Lemma 6, the enlarged radius $\widetilde{\delta}_{n}:=\delta_{n}\vee\{\log(en)/n\}^{1/2}$ satisfies the critical inequality $\mathcal{J}_{2}(\widetilde{\delta}_{n},\mathrm{star}(\overline{\mathcal{F}}_{\ell}))\lesssim\sqrt{n}\,\widetilde{\delta}_{n}^{2}$, up to constants depending on $M$. Combining Theorem 3 with Lemma 5 and Lemma 6 yields the following corollary.
@@
The corollary is stated in terms of the entropy-based critical radius $\delta_{n}$ of the original ERM class $\mathcal{F}$, rather than the critical radius of $\mathrm{star}(\overline{\mathcal{F}}_{\ell})$. This simplification comes at a cost: by Lemma 6, it introduces an additional $\log(eMn)/n$ term in the regret bound. This term is typically loose and can often be improved with more refined arguments; see, for example, Theorem 14.20(b) of Wainwright (2019), Appendix K of Foster and Syrgkanis (2023); Chapter 8 of Sen (2022). In particular, Appendix B.3 gives a local concentration inequality, derived from Theorem 14.20 of Wainwright (2019), that may be preferred to Theorem 2 when the loss $\ell$ satisfies Condition A3. When $\mathcal{F}$ is convex, it replaces $\log(eMn)$ by $\log\log(eMn)$, and it allows the entropy integral to be replaced by the localized Rademacher complexity of $\mathcal{F}$. Alternatively, if one only seeks the in-probability rate $R(\hat{f}_{n})-R(f_{0})=O_{p}(\delta_{n}^{2})$, it suffices to combine Theorem 3.4.1 of Van Der Vaart and Wellner (1996) with Theorem 6 applied directly to $\mathcal{F}_{\ell}$, thereby avoiding star-shaped hulls and the resulting extra logarithmic factors. Nonetheless, the corollary suffices for most nonparametric applications. Outside low-dimensional parametric settings—in which $\delta_{n}^{2}$ can scale as $n^{-1}$ and $\log n/n$ may dominate—it yields the usual sharp rates.
4.3 Illustrative examples
@@ 5.1 Weighted ERM and regret transfer
The following theorem characterizes the excess regret incurred by estimating the weight function. In what follows, let $f_{0}\in\arg\min_{f\in\mathcal{F}}R(f;w_{0})$ denote a population risk minimizer, and let $\hat{f}_{n}$ be any random element of $\mathcal{F}$. For any weight function $w$, define the regret under weight $w$ by
$\mathrm{Reg}(f;w):=R(f;w)-\inf_{f\in\mathcal{F}}R(f;w),$
where $R(f;w):=P\{w(\cdot)\,\ell(\cdot,f)\}.$ We also impose a weighted variant of the Bernstein condition.
@@ 1. B1)
[…]
Remark 1 (Weighted losses are orthogonal under realizability).
Consider the regression setting in which $R(f;w_{0})=E_{P}\{w_{0}(X)\,\ell((X,Y),f)\}$ and $\mathcal{F}$ consists of functions $f(x)$. Appendix G.2 of Foster and Syrgkanis (2023) shows that, under a global realizability condition—typically requiring that the global minimizer lies in $\mathcal{F}$—the weighted loss is orthogonal. Consequently, the quadratic remainder term $\bigl\|1-\tfrac{\hat{w}}{w_{0}}\bigr\|^{2}$ in Theorem 8 can be replaced by a higher-order term, such as $\bigl\|1-\tfrac{\hat{w}}{w_{0}}\bigr\|_{L^{4}(P)}^{4}$. Intuitively, this robustness arises because, when the global minimizer is realizable within $\mathcal{F}$, the choice of weights does not change the population minimizer. As a result, small estimation errors in the weights affect the population risk only at higher order.
@@ 2. B3)
Condition B2 holds under the sufficient conditions of Lemma 1. Condition B3B3-ii ensures that $f$ and $g$ enter the loss $\ell_{g}(\cdot,f)$ in a pointwise Lipschitz manner. The product structure in Condition B3B3-i is satisfied by many losses. For example, it holds for pseudo-outcome regression losses (Rubin and van der Laan, 2006; Yang et al., 2023), such as the DR-learner (van der Laan, 2013; Luedtke and van der Laan, 2016; Kennedy, 2023), which take the form $\ell_{g}(z,f)=\{g(z)-f(z)\}^{2}$ for a pseudo-outcome function $g$. In this case, one may set $m_{1}(z,f):=f(z)$, $m_{2}(z,g):=-2g(z)$, $r_{1}(z,f):=f(z)^{2}$, and $r_{2}(z,g):=g(z)^{2}$. The condition also holds for the weighted loss in Section 5.1, $\ell_{g}(z,f):=g(z)\,\ell(z,f)$, with $m_{1}(z,f):=\ell(z,f)$ and $m_{2}(z,g):=g(z)$. More generally, our analysis extends to losses that can be expressed as finite sums of terms with this product structure, potentially involving multiple nuisance components, including the loss class studied by van der Laan et al. (2024a).
@@ 1. B4)
    ($L^{2}$-to-$L^{\infty}$ interpolation inequality) There exist $c_{\infty}\in(0,\infty)$ and $\beta\ \geq 1/2$ such that
@@ 2. B5)
    (PAC-style error bound for the nuisance) For all $\eta\in(0,1)$,
@@
Condition B4 always holds with $\beta=1/2$ and $c_{\infty}:=2\sup_{f\in\mathcal{F}}\|f\|_{\infty}$, but for smoothness classes it often holds with $\beta>1/2$. For finite-dimensional linear models of dimension $p$, one may take $c_{\infty}\asymp\sqrt{p}$ and $\beta\rightarrow\infty$. It also holds for reproducing kernel Hilbert space balls with eigenvalue decay $\lambda_{j}\asymp j^{-2\beta}$ (Mendelson and Neeman, 2010, Lemma 5.1), for $\ell^{1}$-balls (signed convex hulls) in suitable high- or infinite-dimensional linear spaces (van de Geer, 2014, Lemma 2.1), and for $d$-variate Hölder and Sobolev classes of order $s>d/2$ on bounded domains, where $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). Condition B5 requires that the nuisance estimator satisfy a high-level PAC-style guarantee. For example, this follows from Theorem 3 when the nuisance is itself obtained by empirical risk minimization over $\mathcal{G}$, with $\varepsilon_{\mathrm{nuis}}(n,\eta)$ scaling with the critical radius of $\mathcal{G}$.
The following theorem characterizes the $L^{2}(P)$ error of $\hat{f}_{n}$ in terms of the critical radius of the optimization class $\mathcal{F}$ and the critical radius of the nuisance class $\mathcal{G}$. Its proof relies on a local concentration bound for empirical inner products of the form $\{(P_{n}-P)(fg):\ f\in\mathcal{F},\ g\in\mathcal{G}\}$ (Theorem 13 in Appendix C) and on the localized Rademacher complexity $\mathfrak{R}_{n}\bigl(\mathcal{G},\varepsilon_{\mathrm{nuis}}(n,\eta)\bigr)$ of $\mathcal{G}$. We write $\mathcal{F}$-$\mathcal{F}$ for the pairwise difference class $\{f-f^{\prime}:f,f^{\prime}\in\mathcal{F}\}$, and similarly for $\mathcal{G}-\mathcal{G}$.
@@
$\delta_{n,\mathcal{F}}\wedge\delta_{n,\mathcal{G}}\ \gtrsim\ M\sqrt{\frac{\log(1/\eta)+\log\log(eMn)}{n}}.$
where $M:=1\vee\sup_{g\in\mathcal{G}}\|g\|_{\infty}\vee\sup_{f\in\mathcal{F}}\|f\|_{\infty}$. Then, with probability at least $1-\eta$,
@@
As a special case, taking $\hat{g}=g_{0}$ and $\mathcal{G}:=\{g_{0}\}$, our proof permits setting $\delta_{n,\mathcal{G}}=0$, thereby recovering Theorem 14.20 in Wainwright (2019) for the regret under fixed (known) losses: $\|\hat{f}_{n}-f_{0}\|\lesssim\delta_{n,\mathcal{F}}$. In contrast to Theorem 4, Theorem 9 leverages B3B3-ii to express the regret bound in terms of the critical radius of $\mathcal{F}$-$\mathcal{F}$, rather than that of the loss-difference class.
For simplicity, suppose that $\hat{g}$ is obtained by empirical risk minimization over $\mathcal{G}$. In this case, Theorem 3 typically yields the PAC-style bound $\varepsilon_{\mathrm{nuis}}(n,\eta)\asymp\delta_{n,\mathcal{G}}$, where $\delta_{n,\mathcal{G}}$ is the critical radius of $\mathcal{G}$. Under this specialization, Theorem 9 makes explicit how the critical radii of $\mathcal{F}$ and $\mathcal{G}$ jointly control the $L^{2}(P)$ error of $\hat{f}_{n}$. The leading term, $\delta_{n,\mathcal{F}}^{2}$, matches the oracle rate that would be obtained if the nuisance were known. The additional price of estimating the nuisance in sample appears through $\{\delta_{n,\mathcal{G}}^{2}\}^{4\beta/(2\beta+1)}$. The Hölder exponent $\beta$ in Condition B4 quantifies how effectively $L^{2}(P)$ localization translates into sup-norm control, and therefore governs how strongly the nuisance terms are attenuated. More generally, the theorem highlights two regimes. In the oracle regime, if $\delta_{n,\mathcal{G}}=O(\delta_{n,\mathcal{F}}^{(2\beta+1)/(4\beta)}),$ then the nuisance terms are $O(\delta_{n,\mathcal{F}}^{2})$, and the overall rate matches that of an oracle that knows $g_{0}$; in particular, $\|\hat{f}_{n}-\hat{f}_{0}\|^{2}\lesssim\delta_{n,\mathcal{F}}^{2}$ up to constants. By contrast, if $\delta_{n,\eta,\mathcal{G}}$ exceeds $\delta_{n,\mathcal{F}}^{(2\beta+1)/(4\beta)}$, then the nuisance term dominates $\delta_{n,\mathcal{F}}^{2}$, and the rate is governed by the complexity of the nuisance class.
The following corollary shows that the oracle rate can be attained under the mild Donsker-type requirement that $\delta_{n,\mathcal{G}}^{2}=O(n^{-1/2})$ (and likewise $\varepsilon_{\mathrm{nuis}}^{2}(n,\eta)=O(n^{-1/2})$) for optimization classes satisfying the critical-radius scaling $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$.
@@
When $\delta_{n,\mathcal{G}}\asymp\varepsilon_{\mathrm{nuis}}(n,\eta)$, Corollary 4 yields a simple sufficient condition for oracle performance without sample splitting. Specifically, the oracle rate holds if (i) $\delta_{n,\mathcal{G}}^{2}=O(n^{-1/2})$ and (ii) there exists $\beta>1/2$ such that $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$ and the sup-norm bound in Condition B4 holds with the same $\beta$. Condition (i), equivalently $\delta_{n,\mathcal{G}}=O(n^{-1/4})$, is a Donsker-type requirement on the nuisance class. For example, this condition holds whenever the global Rademacher complexity satisfies $\mathfrak{R}_{n}(\infty,\mathrm{star}(\mathcal{G}-\mathcal{G}))=O(n^{-1/2})$, or, more classically, whenever $\mathcal{J}_{2}(\infty,\mathcal{G})<\infty$, which is sufficient for $\mathcal{G}$ to be Donsker (Van Der Vaart and Wellner, 1996). In turn, the scaling in (ii) is implied if the local Rademacher complexity [^2] The scaling $\sqrt{n}\,\mathfrak{R}_{n}(\delta,\mathcal{F}-\mathcal{F})\asymp\delta^{1-\tfrac{1}{2\beta}}$ need only hold for $\delta\gtrsim\delta_{n,\mathcal{F}}$ to ensure that $\delta_{n,\mathcal{F}}\asymp n^{-\beta/(2\beta+1)}$. and a local $L^{2}(P)$-to-$L^{\infty}$ bound both hold at the same scale:
@@
Requirement (ii) (and (11)) holds for several standard function classes, including Hölder and Sobolev classes, as well as certain RKHS classes. It typically holds for finite-dimensional linear spaces of dimension $k(n)$, with scaling proportional to $\sqrt{k(n)}\,\delta$ (van de Geer, 2014). Theorem 2.3 of van de Geer (2014) further suggests that this scaling is natural for $\ell^{1}$-balls in potentially high- or infinite-dimensional linear spaces. For Hölder and Sobolev classes on $\mathbb{R}^{d}$ with smoothness $s>d/2$, one typically has (11) with $\beta=s/d$ (see Bibaut et al. 2021, Lemma 4 and Adams and Fournier 2003; Triebel 2006). The same scaling holds for a broad class of reproducing kernel Hilbert space balls (RKHSs) whose kernel-operator eigenvalues satisfy $\lambda_{j}\asymp j^{-2\beta}$. In particular, the sup-norm bound in Condition B4 follows from Lemma 5.1 of Mendelson and Neeman (2010), whereas the corresponding critical-radius scaling is established in Theorem 4.7 of Mendelson and Neeman (2010) and Chapter 13 of Wainwright (2019).
The $n^{-1/4}$-rate condition $\varepsilon_{\mathrm{nuis}}(n,\eta)=O(n^{-1/4})$ imposed by (i), together with the scalings in (ii), also appears in the R-learner analyses of Nie and Wager (2021), which study ERM of the CATE with cross-fitted nuisance estimation over RKHSs (see also Example 1 on p. 19 of Foster and Syrgkanis (2023)). This suggests that, at least for RKHSs, the requirements for oracle rates are similar with and without cross-fitting, provided that $\mathcal{G}$ satisfies a Donsker-type condition and that $\delta_{n,\mathcal{G}}\asymp\varepsilon_{\mathrm{nuis}}(n,\eta)$, as is typical when $\hat{g}$ is obtained by ERM over $\mathcal{G}$. Finally, our analysis establishes the oracle rate only up to multiplicative constants; a sharper argument, as in Theorem 5 of van der Laan et al. (2024a), also recovers the oracle constant.